OpenStudy (anonymous):

Here is my last one if you don't mind: (2^x^2+1)/(2^x-1)=16

OpenStudy (anonymous):

I thought 2^x^2?

OpenStudy (anonymous):

I don't know how to write the first part of the exponent. The exponent is x^2 +1

OpenStudy (anonymous):

2^x=t t^2+1=16t-16 now solve it and tell me t's value

OpenStudy (anonymous):

Oh wait, I think I don't understand the question. Is it \(\Large\frac{2^{x^2+1}}{(2^{x-1})}=16\) \[\]

OpenStudy (anonymous):

Yes Ishaan that is correct

OpenStudy (anonymous):

\[2^{x^2 +1 - x +1} = 2^4\] \[x^2 -x + 2=4\]

OpenStudy (anonymous):

Wow You make that look so easy!!! Thanks so much

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2+-x+%2B+2%3D4 At x=2,-1 You're Welcome!

OpenStudy (anonymous):

I have one more question, but I'm going to put it in the Question Panel.

OpenStudy (anonymous):

\[8^{x ^{2}-2x}=0.5\]